SourMunchkin7806:

125(cos 288° + i sin 288°).

SourMunchkin7806:

@Hero @Vocaloid

SourMunchkin7806:

@Ultrilliam

SourMunchkin7806:

@Hero

SourMunchkin7806:

@dude

dude:

You want steps? or are you just lost?

kittybasil:

Hero typically only answers 1 question per month, js

kittybasil:

Also what did you want to do with this?\[\huge{125(\cos{288°}+i\sin{288°})}\]

dude:

Solve it I assume kitty XD

kittybasil:

But solve for what jay? ಠ_ಠ

dude:

J solve, like 2-2=0

jasonmitchell:

=(125*(cos288+isin288))^(1/3) =5*(cos288+isin288)^(1/3) and we know that cosx+isinx can be writeen as e^i(x) , thus : =5*(e^(i288))^(1/3) =5*e^(i*288*1/3) = 5*e^(i96) =5*(cos96+isin96)